A rectangular playground 40m by 60m is to be doubled in area by extending the length and width by an equal amount. By how much should the length and the width be lengthened?
Answer by ankor@dixie-net.com(22740)

A rectangular playground 40m by 60m is to be doubled in area by extending the
length and width by an equal amount.
By how much should the length and the width be lengthened?
:
Find the area of the present playground
40 * 60 = 2400 sq/m
then
2*2400 = 4800 sq/m is the new area
:
Let x = amt to extend the length and the width
(x+40)*(x+60) = 4800
FOIL
x^2 + 60x + 40x + 2400 = 4800
A quadratic equation
x^2 + 100x + 2400 - 4800 = 0
x^2 + 100x - 2400 = 0
Factors to
(x+120)(x-20) = 0
The positive solution
x = 20 m, to extend dimensions to double the area
;
:
See if that's true, find the area with the new dimensions
60 * 80 = 4800